Judy and Jen both drive the same distance to the shore. By driving 15 miles per hour faster than Jen, Judy can get there in 3 hours while Jen takes 4 hours. How fast does each drive?
I'd appreciate help putting this into an equation, which I can then solve.
Thanks!

Answer by edjones(8007)

distance=speed*time
d=s*t
Let x=Jen's speed; x+15= judy's speed
The distance they are travelling is the same so Jan's s*t=Judy's s*t
4x=3(x+15)
4x=3x+45
subtract 3x from each side. x=45 mph
x+15=60 mph
Check:
45*4=180 mi
60*3=180 mi
Ed